OpenStudy (anonymous):

A straight 1.2 m length iron rod is moved at 10 m/s through the earth's magnetic field at a location where it has magnitude 1.0 × 10^-5 T. What is the magnitude of the induced emf between one end of the wire and the other end? A. 1.2 V B. 1.2 × 10^-4 V C. 1.0 × 10^-4 V D. 120 V

OpenStudy (michele_laino):

I'm pondering...

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

during the motion of that iron rod, an electric field will appear inside that iron rod, since all free electrons inside that rod are underwent to the Lorentz force. Now the motion of such free electrons will terminate when the Lorentz force is balanced with the resultant electrostatic force. The situation is as below: 

OpenStudy (anonymous):

yes:)

OpenStudy (michele_laino):

so we can write: \[\Large evB = e\frac{V}{L}\] where V is the emf force, and v is the velocity of the iron rod after a simplification we get: \[\Large V = vBL\]

OpenStudy (michele_laino):

L is the length of the iron rod

OpenStudy (anonymous):

ok! what do we plug in?

OpenStudy (michele_laino):

here is the next step: \[\Large V = vBL = 10 \times {10^{ - 5}} \times 1.2 = ...volts\]

OpenStudy (anonymous):

ok! we get 1.2E-4? so choice B is our solution? :O

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

ooh yay! thanks:)

OpenStudy (michele_laino):

:)